# Uncertainty of magnetic field in 3D NLTE inversions

### v2

@@ 4 Numerical experiment
To evaluate the performance of the solenoidal inversion, we construct a forward model characterized by a simple thermal and density structure, combined with a relatively intricate magnetic field vector. Using this model, we synthesize the emergent spectra, introduce noise, and assess the inversion’s ability to recover the underlying magnetic field configuration. We use the approach outlined in Sect. 2 to ensure the solenoidality condition and the method described in Sect. 3 to estimate the uncertainty in the inferred magnetic field vector.
@@ 4.1 The forward model
To establish the simplest possible thermal structure for benchmarking, we use a constant-property box representing a prominence-like configuration. In Fig. 1 we show the line-center intensity image in a field of view (FOV) including the underlying solar disk. The computational domain spans $5\times 5\times 5\,\mathrm{Mm}^{3}$, with uniform temperature, $T=10^{4}\,\mathrm{K}$, and atomic number density, $N\,[\mathrm{cm}^{-3}]$ such that $\log N=2.1$. The magnetic field is arbitrary and not intended to replicate any realistic prominence magnetic field topology. To evaluate the method’s capability to infer the magnetic field vector and quantify the associated uncertainties, we have chosen the following vector potential components, [^3] For notation simplicity, we disregard the normalization factor $\mu(A)$ in these expressions.
@@
The model atom chosen in this study is purely academic, represented by a two-level, normal Zeeman triplet line. The critical Hanle field is set to $B_{\rm H}=284\,\mathrm{G}$, with an upper-level Landé factor of $g_{u}=1$, an atomic mass of $2\,m_{\rm u}$, and a line-center wavelength of $5000\,\mathrm{\AA}$. The domain’s line-center optical thickness is $\tau=1.5$. For boundary illumination, we use the center-to-limb intensities at this wavelength by Allen (1976), accounting for the spherical geometry of the underlying Sun.
The synthetic observation consists of a $64\times 64$ pixel grid, with added noise characterized with a standard deviation of $\sigma=2\times 10^{-9}\,\mathrm{erg\,cm^{-2}\,s^{-1}\,Hz^{-1}\,sr^{-1}}$, which corresponds to approximately $2\times 10^{-4}$ fraction of the line-center intensity of the prominence. The field of view for this observation is shown in Fig. 3.
@@ 4.2 Inversion
The inversion process begins with an initial guess with randomized values for temperature and density and zero magnetic field. We model the spatial variation of these quantities with a set of basis functions of order 4 along all Cartesian axes. Randomized angular quadratures from Štěpán et al. (2020) of order $L=15$ are used, following the method described in Štěpán et al. (2022). The solution was obtained in approximately 16.5 kh of CPU time via a massively parallel run of POLARIS. The final inversion yields a value of $\chi^{2}=1.03$.
@@
In Fig. 4, we show correlation plots, comparing the Cartesian components of the magnetic field in the inverse and true models at $10^{5}$ randomly chosen points in the model’s domain. Despite the relatively small optical thickness of the medium (the smaller the optical thickness, the less significant the expected radiative coupling between different regions), the Pearson correlation coefficients for all components exceed 0.9, indicating a significant accuracy of the inversion.
Following the strategy outlined in Sect. 3, we use an MC approach to estimate the uncertainties of the inferred magnetic field. We generate 100 magnetic field models, each with a randomized vector potential. The randomization is achieved by perturbing all the basis function coefficients, with a variance of $\sigma^{2}=0.3^{2}$ in the units of the normalization factor $\mu(A)$. All 100 inversions converge to a $\chi^{2}$ value below 1.5, and thus all of them are included in the subsequent analysis.
Next, we analyze the uncertainties in the magnetic field inversion. Figure 5 shows histograms of the magnetic field components resulting from the ensemble of randomized models at a specific coordinate in the domain. The peaks of the distributions closely align with the true values of the magnetic field component, something that holds true for most regions of the domain.
@@ 5 Discussion and conclusions
In this work, we have presented a method to ensure the solenoidality of the magnetic field vector inferred in 3D inversions, as well as to provide an estimation of the uncertainty of the inverted magnetic field. One important consideration when using the method is the degree of smoothness in the inferred magnetic fields. While this constraint may prevent the retrieval of small-scale turbulence, we think it is preferable to obtain a globally smooth solenoidal field rather than an arbitrary local field with uncertain solenoidality.
We have found that even a modest optical thickness of about 1.5 along the LOS is sufficient to provide high-quality magnetic field information at a given noise level. This finding is consistent with the results by Štěpán et al. (2022), which emphasized the importance of optical thickness for the accuracy of magnetic field retrievals in the self-consistent models.
@@
Although we did not encounter any ambiguous solutions in our thermodynamically simplest possible model, the issue of ambiguities in the magnetic field inference remains a key topic for future research. This will include the theoretical investigation of the role of noise in the observation and the plasma’s optical thickness in generating such ambiguities. Additionally, the uncertainties in the thermodynamic quantities, which were not addressed in this work, warrant further investigation to assess their impact on the inversion process.

### v3

@@ 4 Numerical experiment
To evaluate the performance of the solenoidal inversion, we constructed a forward model characterized by a simple thermal and density structure, combined with a relatively intricate magnetic field vector. Using this model, we synthesized the emergent spectra, introduce noise, and assess the inversion’s ability to recover the underlying magnetic field configuration. We used the approach outlined in Sect. 2 to ensure the solenoidality condition and the method described in Sect. 3 to estimate the uncertainty in the inferred magnetic field vector.
@@ 4.1 The forward model
To establish the simplest possible thermal structure for benchmarking, we used a constant-property box representing a prominence-like configuration. In Fig. 1 we show the line-center intensity image in a FOV including the underlying solar disk. The computational domain spans $5\times 5\times 5\,\mathrm{Mm}^{3}$, with uniform temperature, $T=10^{4}\,\mathrm{K}$, and atomic number density, $N\,[\mathrm{cm}^{-3}]$ such that $\log N=2.1$. The magnetic field is arbitrary and not intended to replicate any realistic prominence magnetic field topology. To evaluate the method’s capability to infer the magnetic field vector and quantify the associated uncertainties, we chose the following vector potential components, [^3] For notation simplicity, we disregard the normalization factor $\mu(A)$ in these expressions.
@@
The model atom chosen in this study is purely academic, represented by a two-level, normal Zeeman triplet line. The critical Hanle field is set to $B_{\rm H}=284\,\mathrm{G}$, with an upper-level Landé factor of $g_{u}=1$, an atomic mass of $2\,m_{\rm u}$, and a line-center wavelength of $5000\,\mathrm{\AA}$. The domain’s line-center optical thickness is $\tau=1.5$. For boundary illumination, we used the center-to-limb intensities at this wavelength from Allen (1976), which account for the spherical geometry of the underlying Sun.
The synthetic observation consists of a $64\times 64$ pixel grid, with added noise characterized with a standard deviation of $\sigma=2\times 10^{-9}\,\mathrm{erg\,cm^{-2}\,s^{-1}\,Hz^{-1}\,sr^{-1}}$, which corresponds to approximately $2\times 10^{-4}$ fraction of the line-center intensity of the prominence. The FOV for this observation is shown in Fig. 3.
@@ 4.2 Inversion
The inversion process begins with an initial guess with randomized values for temperature and density and zero magnetic field. We modeled the spatial variation of these quantities with a set of basis functions of order 4 along all Cartesian axes. Randomized angular quadratures from Štěpán et al. (2020) of order $L=15$ are used, following the method described in Štěpán et al. (2022). The solution was obtained in approximately 16.5 kh of CPU time via a massively parallel run of POLARIS. The final inversion yields a value of $\chi^{2}=1.03$.
@@
In Fig. 4 we show correlation plots that compare the Cartesian components of the magnetic field in the inverse and true models at $10^{5}$ randomly chosen points in the model’s domain. Despite the small optical thinness of the medium (the smaller the optical thickness, the less significant the expected radiative coupling between different regions), the Pearson correlation coefficients for all components exceed 0.9, indicating a significant accuracy of the inversion.
Following the strategy outlined in Sect. 3, we used an MC approach to estimate the uncertainties of the inferred magnetic field. We generated 100 magnetic field models, each with a randomized vector potential. The randomization is achieved by perturbing all the basis function coefficients, with a variance of $\sigma^{2}=0.3^{2}$ in the units of the normalization factor $\mu(A)$. All 100 inversions converge to a $\chi^{2}$ value below 1.5, and thus all of them are included in the subsequent analysis.
Next, we analyzed the uncertainties in the magnetic field inversion. Figure 5 shows histograms of the magnetic field components resulting from the ensemble of randomized models at a specific coordinate in the domain. The peaks of the distributions closely align with the true values of the magnetic field component, something that holds true for most regions of the domain.
@@ 5 Discussion and conclusions
We have presented a method for ensuring the solenoidality of the magnetic field vector inferred in 3D inversions, as well as for providing an estimation of the uncertainty of the inverted magnetic field. One important consideration when using the method is the degree of smoothness in the inferred magnetic fields. While this constraint may prevent the retrieval of small-scale turbulence, we think it is preferable to obtain a globally smooth solenoidal field rather than an arbitrary local field with uncertain solenoidality.
We have found that even a modest optical thickness of about 1.5 along the LOS is sufficient to provide high-quality magnetic field information at a given noise level. This finding is consistent with the results by Štěpán et al. (2022), which emphasize the importance of optical thickness for the accuracy of magnetic field retrievals in self-consistent models.
@@
Although we did not encounter any ambiguous solutions in our thermodynamically simplest possible model, the issue of ambiguities in the magnetic field inference remains a key topic for future research. This will include a theoretical investigation of the role of noise in the observation and the plasma’s optical thickness in generating such ambiguities. Additionally, the uncertainties in the thermodynamic quantities, which were not addressed in this work, warrant further investigation to assess their impact on the inversion process.